TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-19 23:26:00
subject: 3\10 Pt-1 ISS On Orbit Status 10-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\10 ISS On-Orbit Status 10-03-2003
Part 1 of 2

ISS On-Orbit Status 3/10/03

All ISS systems continue to function nominally, except as noted 
previously or below.

Week 15 for Expedition 6 started for FE-2/SO Don Pettit with continued 
MSG (Microgravity Science Glovebox) troubleshooting. Status: So far, 
so good.  [Pettit reconnected electronic box connector P105, which 
powers the Work Volume side loading port sensors and the CMP test 
interface panel. If ESEM3 (exchangeable standard electronic module 3) 
current remained nominal, connector P103, for the left, center, and 
right fan inverters, was to be re-installed next today, in the first 
of several back-to-back runs slated for this week. Early analysis has 
shown no off-nominal current draw on ESEM3 as a result of the P106 and 
P101 connectors re-installed this weekend. At this phase of 
troubleshooting, the MSG should slowly begin to show more "visual 
cues" such as additional H&S LEDs (health and status light-emitting 
diodes) illuminating.]

FE-1 Nikolai performed the periodic air sampling in the station, first 
using the standard Russian AK-1M sampler device in the Service Module 
(SM), FGB and Lab. Then, to test the SM for Freon, he used the AK-1M-F 
air sampler.  Later, testing for CO (carbon monoxide) levels, he 
collected air samples in the SM and DC-1 docking compartment with the 
IPD Draeger tubes sampler.

CDR Ken Bowersox completed another pre-EVA EVARM (EVA radiation 
monitoring) background reading, first activating the EVARM badge 
reader, then taking the readings. Afterwards the reader was powered 
down again.

In preparation for the Stage EVA2 by Bowersox and Pettit (tentatively
scheduled for 3/25), Don Pettit initiated recharge of the first two 
(of four) EMU (extravehicular mobility unit) batteries in the BSA 
(battery stowage assembly). [A consumables plan for the EVA was 
uplinked overnight, which, besides battery management, also specifies 
the recyclable Metox (metal oxide) canisters to be used, as well as 
the ground-recommended SAFERs (simplified aid for EVA rescue, ##1003 & 
1004) and a revised cooling water recharge procedure.]

Also uplinked were updated instructions for using the onboard SAFER 
trainer in an SSC (station support computer)-based onboard training 
run placed on the U.S. task list for this week.

Later in the day, in a teleconference with ground specialists, the 
crew was treated to a technical "refresher" course on fluid QDs (quick 
disconnects) and their corrective SPDs (spool positioning devices) for 
the spacewalk.

Budarin performed his daily status check on the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment in the SM.

Nikolai also completed the daily routine maintenance of the SM's SOSh 
life support system, including ASU toilet facilities, the preparation 
of the IMS inventory delta file for downlink, and the regular 
inspection of the BRPK air/condensate separator of the SRVK water 
processing system. Early in the morning, Bowersox conducted the 
regular status checkup of the autonomous PCG-STES010 payload in the 
Lab.

All crewmembers performed their regular daily physical exercise on 
TVIS (treadmill with vibration isolation and stabilization), RED 
(resistive exercise device), and VELO ergometer with load trainer.

Uplink/upload procedures for the new software patch to correct the R3
anomaly for the GNC MDMs (guidance, navigation and control computers) 
are being tested at MCC-H today. If the test is successful, 
installation of the patch will take place next Wednesday (3/12) in 
conjunction with the handover of ISS attitude control to the Russian 
segment for the Progress thruster testing.

The test burns of the two Progress 10P thruster manifolds are still 
set for 3/12 (5:23pm EST) and 3/13 (6:38pm) this week, each one a 
mini-reboost of 1.38 m/sec delta-V for the first, 0.31 m/sec for the 
second burn.

Tonight during crew sleep, MCC-Moscow will conduct the video survey of 
the SM's #2 solar array with the portside Kurs-P antenna (AKP) and its
high-frequency feeder cable, which runs along the array's center boom 
and is suspected of being damaged. [The SSRMS (space station remote 
manipulator system) was maneuvered in the proper position last week, 
and MCC-H will have the onboard MSS (mobile service system) powered up 
by at 4:45pm EST. This will allow subsequent camera setup during the 
Ku pass starting at 5:17pm.  The robotarm's Tip LEE (latching end 
effector), Tip Elbow and Base Elbow cameras will be used, routed to 
downlink via Ku-band, with the Kurs-AKP survey starting at 6:16pm for 
about ten minutes (backup 7:55pm), followed by a survey of the SVS 
(space vision system) target on the P1 truss, probably starting at 
6:56pm.]

For troubleshooting the IMV (intra-module ventilation) fan, which the 
crew reported last week as noisy, Bowersox and Pettit worked on the 
Node and Lab IMV air duct system.

MCC-H Flight Surgeon requested OCA comm. downlink of accumulated MEC
(medical equipment computer) files for CEVIS (cycle ergometer with 
vibration isolation), HRM (heart rate monitor), RED, TVIS, and FFQ 
(food frequency questionnaire) nutrition log.

In the interest of conserving precious resources, Budarin completed 
another extensive 3-hr. IMS inventory audit in the FGB "Zarya" module, 
of up to 40 items shown on the IMS database as stowed behind Panel 
227. Audit objective was to update the status of dust 
collector/filters and cartridges as well as U.S. hardware 
availability.

Today's targets for the CEO (crew earth observations program) were 
Western Mediterranean Dust (light was low at the time of this pass, 
but the crew was asked to look right of track for morning, sunlit dust 
over the dark surface of western Med), Addis Ababa (Addis Abeba), 
Ethiopia (few, if any clouds were expected this nadir pass over the 
Ethiopian capital), Casablanca, Morocco (if coastal stratus cloud 
formations remained offshore, ISS/CEO had a nadir pass over this 
famous Moroccan coastal city), Mozambique Flooding (DYNAMIC EVENT 
TARGET: The remains of Tropical Cyclone Japahet lingered over central 
Mozambique for most of this past week dumping torrential rains. Crew 
was to document conditions in the Zambezi and Save River basins 
looking for evidence of lowland flooding, sediment plumes, and 
vegetation contrast from wind and water damage).

CEO images can be viewed at the website
http://eol.jsc.nasa.gov
http://earthobservatory.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.